# SPAMming Labels: Efficient Annotations for the Trackers of Tomorrow

### v1

@@ Abstract
Increasing the annotation efficiency of trajectory annotations from videos has the potential to enable the next generation of data-hungry tracking algorithms to thrive on large-scale datasets. Despite the importance of this task, there are currently very few works exploring how to efficiently label tracking datasets comprehensively. In this work, we introduce SPAM, a tracking data engine that provides high-quality labels with minimal human intervention. SPAM is built around two key insights: i) most tracking scenarios can be easily resolved. To take advantage of this, we utilize a pre-trained model to generate high-quality pseudo-labels, reserving human involvement for a smaller subset of more difficult instances; ii) handling the spatiotemporal dependencies of track annotations across time can be elegantly and efficiently formulated through graphs. Therefore, we use a unified graph formulation to address the annotation of both detections and identity association for tracks across time. Based on these insights, SPAM produces high-quality annotations with a fraction of ground truth labeling cost. We demonstrate that trackers trained on SPAM labels achieve comparable performance to those trained on human annotations while requiring only $3-20\%$ of the human labeling effort. Hence, SPAM paves the way towards highly efficient labeling of large-scale tracking datasets. Our code and models will be available upon acceptance.
@@
This supplementary material offers additional details (Appendix A) and experimental results (Appendix B) of our SPAM video data label engine. To this end, we provide further information on the architecture of our graph-based model (Section A.1) and its application for hierarchical labeling (Section A.2). In Section A.3, we complete the implementation details of the main paper. The experimental results in Section B.1 demonstrate the superiority of our approach to a common frame-based labeling method. Finally, we compare labeling performance for static and moving cameras (Appendix B).
00footnotetext: * Work done during an internship at NVIDIA.
@@ 1 Introduction
Detecting and tracking objects in time is at the core of many challenging vision applications. To address these problems, modern multi-object tracking (MOT) approaches [46, 79, 81, 75, 69] rely on increasingly larger amounts of annotated data. While annotating image datasets is already costly, introducing a temporal component further increases the task difficulty and data scale requirements. This is due to the fact that redundancies between frames cause the information density to scale poorly with the amount of data, making the overall annotation task more challenging and resource-intensive.
@@
To mitigate these issues, researchers are spending more effort in operating with limited human annotations. Self- and semi-supervised learning (SSL) and pseudo-labeling (PL) approaches are often used to leverage unlabeled data whenever possible [36, 42, 35], while Active learning (AL) strategies aim at selecting the most informative samples to label. All these strategies are commonly applied in various image understanding tasks, such as classification [14, 11] or object detection [17, 70]. Even if video tasks do not have any lower requirement for data, there are only a handful of works that target efficient labeling in the video domain [62, 63, 64, 40]. Existing works fail to provide comprehensive solutions to video labeling by either ignoring its dense temporal component [40, 64] or working under a limited single-object setup [74]. An effective MOT labeling engine should operate across multiple objects and frames, considering the intricate dependencies between tracks inherent in the labeling problem.
@@
Videos have unique characteristics that can be leveraged to unleash the full potential of a modern MOT label engine. Firstly, frame-wise similarities in videos yield data redundancies. In the context of MOT, this implies that some associations between detections are relatively easy to solve [7]. This setup provides an ideal opportunity to leverage state-of-the-art models pre-trained to pseudo-label new data at zero cost. Secondly, the dependencies of objects across time and space cause annotations in one frame of the video to have cascading effects in other frames. For example, solving an association for one track can resolve associations for neighboring tracks. In order to maximize the efficiency of the annotations, we frame the labeling task within a model that takes into account such dependencies. Thus, we adopt a track-centric labeling approach rather than a frame-centric one [40].
@@
We present the SPAM video data label engine, which chains Synthetic pre-training, Pseudo-labeling, and Active learning with a graph-based Model for optimal tracking performance with minimal human effort. In the image domain, there are plenty of large-scale annotated datasets that a model-driven approach can leverage for pre-training [35], but this is not the case for multi-object tracking datasets, which tend to be much smaller in scale. To avoid depending on an initial well-curated large-scale dataset, we advocate for a Synthetic pre-training. We further analyze which components of a tracking method are more affected by the synthetic-to-real gap. This first experiment gives us an important insight: labeling should be focused on detections and associations, and we can directly use the synthetically trained reID model. With our synthetic pre-trained model, we then propose a model-driven approach to produce Pseudo-labels on real data and retrain the model with its own pseudo-labels. We show that with this recipe we can even reach a performance comparable to ground truth without ever training our labeler on manually annotated data.
@@
Synthetic pre-training and pseudo-labeling allow us to handle easy decisions and defer only uncertain or complex decisions to an annotator. We propose an Active learning strategy where the labeler makes decisions at the track-level instead of frame-level, thereby utilizing the annotation budget more efficiently. Towards this goal, we solve the labeling problem with a Model, based on graph hierarchies [12]. This model captures long-range spatio-temporal dependencies, which means our annotations will have an impact across frames, and not only on the frame where the annotation actually happens. We further enhance the model with a detection filtering layer so that we can naturally label detection and association within the same framework.
@@
In summary, we propose SPAM, a multi-object tracking data label engine that significantly reduces annotation cost by dealing with easy scenarios with (i) a strong MOT model pre-trained on Synthetic data, and used to (ii) Pseudo-label real data for its own re-training. Harder detection and association decisions are carefully selected by an (iii) Active learning scheme that works on a (iv) Model based on graph hierarchies. This allows us to efficiently use the annotator’s budget and account for spatio-temporal dependencies in the labels. The resulting model generates autolabels comparable to ground truth performance, requiring only $3-20\%$ of manual annotation effort across three diverse datasets.
@@ 2 Related Work
As a comprehensive video data label engine, SPAM is the first of its kind to label tracking datasets by uniquely chaining the following four aspects.
@@
Our label engine also pre-trains the labeling model synthetically, minimizing its total annotation footprint. To this end, we demonstrate effectiveness of synthetic pre-training [23] for generating pseudo-labels for self-supervision.
@@
Efficient labeling of video datasets. While significant effort has been devoted towards increasing the labeling efficiency of image datasets [1, 35], similar approaches for the video and tracking domain are still relatively limited. Earlier works reduce the number of frames to annotate within individual videos but still rely heavily on human annotators [65]. Other solutions are limited to only a single object per sequence [63, 16] or still rely on human bounding box annotations [62] to label segmentation masks across time. The authors of [41] select key frames based on their expected impact on tracking performance. While effective within its scope, this approach only yields sparse frame annotations, which are insufficient for training most modern trackers. To reduce the human annotation effort for trajectories, [45] proposes an efficient video interface that still requires labeling all trajectories.
@@
For our label engine to generate long-term pseudo-labels and active learning uncertainty estimations, we rely on a strong graph-based tracker [12]. To incorporate both detection and association labeling within a unified graph framework, we introduce a novel hierarchy level dedicated to detection filtering.
@@ 3.1 Overview
Figure 1: Overview of the SPAM training and annotation pipeline. (a) Initial model training on synthetic data. (b) Application of SPAM to generate pseudo-labels without incurring manual annotation costs on a real dataset, followed by training on pseudo-labels. (c) Real dataset labeling using pseudo-labels and an uncertainty-based active learning approach.
Our objective is to minimize the manual annotation effort required for multi-object tracking datasets. Our key observation is that most association cases can be solved easily by models pre-trained on synthetic data. Therefore, no human effort is needed, and we can instead use our model’s predictions as pseudo-labels. Decisions that are more complex or uncertain are be selected by an active learning scheme to be annotated manually. The resulting set of labels achieves near-ground truth (GT) performance with a dramatically reduced need for human annotation, thereby vastly enhancing the efficiency of the overall labeling process.
In order to perform pseudo-labeling and active learning, we require a model that can capture long-term spatio-temporal dependencies between tracks. We leverage a hierarchical graph neural network formulation and cast the annotation task over a graph. Our architecture draws inspiration from SUSHI [12], which demonstrated remarkable association performance. Our model’s scope, however, goes beyond association. Given the constraints of the labeling task, we need to equip our model with the ability to reason about detections before performing association. To address this broader goal, our GNN hierarchy, unlike SUSHI [12], also includes the capability to accept or reject detections by classifying nodes into correct and incorrect object hypotheses.
Our overall model is pre-trained on synthetic data [23] and further fine-tuned on its own predicted pseudo-labels over the target dataset that we aim to annotate. This recipe allows for strong labelling performance without any human annotations, and provides the basis for a scalable video annotation engine. After this training loop, we use our updated model’s predictions as pseudo-labels to annotate a large portion of the target dataset. Decisions that are more complex or uncertain are selected by an active learning scheme to be annotated manually. The resulting set of labels achieves near-GT performance with a dramatically reduced set of manual annotations, vastly enhancing the efficiency of the overall process. We provide an overview of our training and annotating recipe in Fig. 1.
@@ 3.2 Graph Hierarchies as a Model for Labeling
Problem formulation. Our model builds upon the standard graph formulation [76] where, given a set of object candidates $\mathcal{O}$, our goal is to obtain a set of valid objects $\mathcal{O}_{v}\subset\mathcal{O}$ and their corresponding trajectories $\mathcal{T}$. Each trajectory $T_{k}\in\mathcal{T}$ consists of a set of objects sharing the same identity or, equivalently, the set of edges connecting those nodes. Specifically, we model each object candidate $o_{i}\in\mathcal{O}$ as a node and each association hypothesis among candidates as an edge in an undirected graph $G=(V,E)$. Within this formulation, nodes $u\in V$ could be classified into valid objects if $u\in\mathcal{O}_{v}$. Analogously, edges can be classified into correct or incorrect hypotheses depending on whether they belong to some trajectory or not. This formulation can be generalized to a hierarchical setting by splitting video sequences into non-overlapping subsequences, i.e., disjoint subgraphs, and progressively merging them. For further details we refer to the supplementary material.
@@
Background on graph hierarchies for MOT. Our model consists of a hierarchy of GNNs operating over the graph formulation described above. Intuitively, starting from detection candidates, at each hierarchy level, GNN hierarchy subsequently merge tracklets from the previous level into longer ones. Nodes and edges are represented by their embeddings. Message passing is then used to share information within the graph in order to update node and edge embeddings with richer information. The GNNs are trained to classify edges into active and inactive association hypotheses and they share the same features, network architecture and learnable weights across hierarchy levels. In addition to the edge features reported in  [12], for each object candidate $o_{i}$ we concatenate its timestamp $t_{i}$, normalized bounding box coordinates and dimensions $(x_{i},y_{i},h_{i},w_{i})$ and corresponding confidence from the detector $c_{i}$ and embed it as node features.
@@
Classifying nodes. We aim to minimize the annotation effort; therefore, we restrict the annotator to operate on a fixed graph of detection candidates i.e. no new detections are added to the graph after the initial set. This is why we have to create an initial graph that is over-complete. In order to do so, we use a low confidence threshold for our detector and obtain a large number of detection candidates (Fig. 2). While this ensures a high detection recall, it also introduces a larger number of false positive detections that our model needs to address. As we show in experiments, the original SUSHI architecture [12] cannot handle such false positives, which is why we propose a new layer in the hierarchy, $GNN_{node}$. Intuitively, $GNN_{node}$ leverages temporal information to validate object candidates. We use its output node embeddings and feed them to a lightweight classifier to enable false positive detection filtering. By doing so, we can consider a large number of input proposals achieving high detection recall by leveraging the spatio-temporal reasoning ability of our GNN for detection candidates.
@@ 3.3 How to Label on Graph Hierarchies?
Figure 3: Our graph-based labeling pipeline begins with the selection of nodes for annotation. For each node to be labeled, the annotator is presented with the options of validating the detection, improving localization by refining the box and association.
Graph-based labeling. Utilizing a graph formulation for detection and tracking enables explicit modeling of objects in the scene and their interactions. In this framework, each node of the graph represents an object, allowing a manual annotation of an object to be framed as selecting a node on this graph. Given a node, we present the annotator with options of: accepting a detection and refining the bounding-box to improve localization (two clicks, one for the top-left corner and one for the bottom-right corner of the bounding box) and associating a node across time (1 click). As we are working on an over-complete set of detections, the annotator can also reject a detection (1 click). We visualize our annotation pipeline in Fig. 3.
Annotation acquisition. The core idea of our approach is to use our model decisions to guide the annotation process. For the vast majority of scenarios, we directly use our model’s predictions as pseudo-labels. For a small subset of more challenging decisions where our model exhibits uncertainty, we validate our decisions with an annotator. Within our graphs, we adopt an uncertainty-based approach to identify which nodes require annotations. Specifically, for each node $v\in V$, we compute its uncertainty as $\text{uncert}(v)\vcentcolon=\text{max}_{u\in\text{N}_{v}}H({\hat{y}_{(v,u)}})$, where $\text{N}_{v}$ denotes the set of neighboring nodes of $v$ and $\hat{y}_{(v,u)}$ is the edge prediction for edge $(v,u)$:
@@
At the $GNN_{node}$ level, where we have node predictions instead of edge predictions, we simply compute $\text{uncert}(v)$ as $H({\hat{y}_{(v)}})$ instead. Subsequently, we refer the nodes with the highest uncertainty scores to the annotator for manual annotation and use the remaining predictions as pseudolabels for the rest of the nodes.

### v2

@@ Abstract
Increasing the annotation efficiency of trajectory annotations from videos has the potential to enable the next generation of data-hungry tracking algorithms to thrive on large-scale datasets. Despite the importance of this task, there are currently very few works exploring how to efficiently label tracking datasets comprehensively. In this work, we introduce SPAM, a video label engine that provides high-quality labels with minimal human intervention. SPAM is built around two key insights: i) most tracking scenarios can be easily resolved. To take advantage of this, we utilize a pre-trained model to generate high-quality pseudo-labels, reserving human involvement for a smaller subset of more difficult instances; ii) handling the spatiotemporal dependencies of track annotations across time can be elegantly and efficiently formulated through graphs. Therefore, we use a unified graph formulation to address the annotation of both detections and identity association for tracks across time. Based on these insights, SPAM produces high-quality annotations with a fraction of ground truth labeling cost. We demonstrate that trackers trained on SPAM labels achieve comparable performance to those trained on human annotations while requiring only $3-20\%$ of the human labeling effort. Hence, SPAM paves the way towards highly efficient labeling of large-scale tracking datasets. We release all models and code.
@@
This supplementary material offers additional details (Appendix 0.A) and experimental results (Appendix 0.B) of SPAM. To this end, we provide further information on the architecture of our graph-based model (Sec. 0.A.1) and its application for hierarchical labeling (Sec. 0.A.2). In Sec. 0.A.3, we complete the implementation details of the main paper. The experimental results in Sec. 0.B.1 demonstrate the superiority of our approach to a common frame-based labeling method. We compare labeling performance for static and moving cameras in Sec. 0.B.2 and finally discuss about an alternative evaluation setting in Sec. 0.B.3.
Keywords:
Multi-object tracking Data labeling Active learning
00footnotetext: $\dagger$ Work done during an internship at NVIDIA.
@@ 1 Introduction
Detecting and tracking objects in time is at the core of many challenging vision applications. To address these problems, modern multi-object tracking (MOT) approaches [45, 78, 80, 74, 68] rely on increasingly larger amounts of annotated data. While annotating image datasets is already costly, temporal component of videos further increases the annotation difficulty and data scale requirements. This is due to the fact that redundancies between frames cause the information density to scale poorly with the amount of data, making the overall annotation task more challenging and resource-intensive.
@@
To mitigate these issues, researchers are spending more effort in operating with limited human annotations. Self- and semi-supervised learning (SSL) and pseudo-labeling (PL) approaches are often used to leverage unlabeled data whenever possible [36, 41, 35], while active learning (AL) strategies aim at selecting the most informative samples to label. All these strategies are commonly applied in various image understanding tasks, such as classification [14, 11] or object detection [17, 69]. Even if video tasks do not have any lower requirement for data, there are only a handful of works that target efficient labeling of tracking data in the video domain [61, 62, 63, 40]. Existing works fail to provide comprehensive solutions to labeling by either ignoring dense temporal component of videos [40, 63] or working under a limited single-object setup [73]. An effective video labeling engine should operate across multiple objects and frames, considering the intricate dependencies between tracks inherent in the labeling problem.
@@
Videos have unique characteristics that can be leveraged to unleash the full potential of a modern label engine. Firstly, frame-wise similarities in videos yield data redundancies. In the context of MOT, this implies that some associations between detections are relatively easy to solve [7]. This setup provides an ideal opportunity to leverage state-of-the-art models pre-trained to pseudo-label new data at zero cost. Secondly, the dependencies of objects across time and space cause annotations in one frame of the video to have cascading effects in other frames. For example, solving an association for one track can resolve associations for neighboring tracks. In order to maximize the efficiency of the annotations, we frame the labeling task within a model that takes into account such dependencies. Thus, we adopt a track-centric labeling approach rather than a frame-centric one [40].
@@
We present the SPAM video label engine, which chains Synthetic pre-training, Pseudo-labeling, and Active learning with a graph-based Model for optimal tracking performance with minimal human effort. In the image domain, there are plenty of large-scale annotated datasets that a model-driven approach can leverage for pre-training [35]. However, this is not the case for multi-object tracking datasets, which are typically much smaller in scale; therefore, we advocate for a Synthetic pre-training. We further analyze which components of a tracking method are more affected by the synthetic-to-real gap. This first experiment gives us an important insight: labeling should be focused on detections and associations, and we can directly use the synthetically trained reID model. With our synthetic pre-trained model, we then propose a model-driven approach to produce Pseudo-labels on real data and retrain the model with its own pseudo-labels. We show that with this recipe we can even reach a performance comparable to ground truth without ever training our labeler on manually annotated data.
@@
Synthetic pre-training and pseudo-labeling allow us to handle easy decisions and defer only uncertain or complex decisions to an annotator. We propose an Active learning strategy where the labeler makes decisions at the track-level instead of frame-level, thereby utilizing the annotation budget more efficiently. Towards this goal, we solve the labeling problem with a Model, based on graph hierarchies [12]. This model captures long-range spatio-temporal dependencies, which means our annotations will have an impact across frames, and not only on the frame where the annotation actually happens. We extend the model with a detection filtering layer so that we can naturally label detection and association within the same framework.
@@
In summary, we propose SPAM, a video label engine that significantly reduces annotation cost by dealing with easy scenarios with a strong MOT model pre-trained on (i) Synthetic data, followed by (ii) Pseudo-labeling real data for its own re-training. Harder detection and association decisions are carefully selected by an (iii) Active learning scheme that works on a (iv) Model based on graph hierarchies. This allows us to efficiently use the annotator’s budget and account for spatio-temporal dependencies in the labels. The resulting model generates labels comparable to ground truth performance, requiring only $3-20\%$ of manual annotation effort across three diverse datasets.
@@ 2 Related Work
Our label engine also pre-trains the labeling model synthetically, minimizing its total annotation footprint. To this end, we demonstrate the effectiveness of synthetic pre-training [23] for generating pseudo-labels for self-supervision.
@@
Efficient labeling of video datasets. While significant effort has been devoted towards increasing the labeling efficiency of image datasets [1, 35], similar approaches for the video and tracking domain are still relatively limited. Earlier works reduce the number of frames to annotate within individual videos but still rely heavily on human annotators [64]. Other solutions are limited to only a single object per sequence [62, 16] or still rely on human bounding box annotations [61] to label segmentation masks across time. The authors of [40] select key frames based on their expected impact on tracking performance. While effective within its scope, this approach only yields sparse frame annotations, which are insufficient for training all modern trackers. To reduce the human annotation effort for trajectories, [44] proposes an efficient video interface that still requires labeling all trajectories.
@@
For our label engine to generate long-term pseudo-labels and uncertainty estimations, we rely on a strong graph-based tracker [12]. To incorporate both detection and association labeling within a unified graph framework, we introduce a novel hierarchy level dedicated to detection filtering.
@@ 3.1 Overview
Our main goal is to label multi-object tracking datasets with minimal annotation effort while enabling optimal tracking performance. Our key observation is that most association cases can be solved easily by models pre-trained on synthetic data. Therefore, no human effort is needed, and we can instead use our model’s predictions as pseudo-labels. Decisions that are more complex or uncertain are selected by an active learning scheme to be annotated manually.
In order to perform pseudo-labeling and active learning, we require a model that can capture long-term spatio-temporal dependencies between tracks. We leverage a hierarchical graph neural network formulation and cast the annotation task over a graph. Our architecture (Fig. 1) draws inspiration from SUSHI [12], which demonstrated remarkable association performance. Our model’s scope, however, goes beyond association. Given the constraints of the labeling task, we need to equip our model with the ability to reason about detections before performing association. To address this broader goal, our GNN hierarchy, also includes the capability to accept or reject detections by classifying nodes into correct and incorrect object hypotheses.
Our overall model is pre-trained on synthetic data [23] and further fine-tuned on its own predicted pseudo-labels over the target dataset that we aim to annotate. This recipe allows for strong labelling performance without any human annotations, and provides the basis for a scalable video label engine. After this training loop, we use our updated model’s predictions as pseudo-labels to annotate a large portion of the target dataset. Decisions that are more complex or uncertain are selected by an active learning scheme to be annotated manually. The resulting set of labels achieves near-GT performance with significantly fewer manual annotations, thereby vastly enhancing the efficiency of the overall process. We provide an overview of our training and annotating recipe in Fig. 2.
@@ 3.2 Graph Hierarchies as a Model for Labeling
Problem formulation. Our model builds upon the standard graph formulation [75] where, given a set of object candidates $\mathcal{O}$, our goal is to obtain a set of valid objects $\mathcal{O}_{v}\subset\mathcal{O}$ and their corresponding trajectories $\mathcal{T}$. Each trajectory $T_{k}\in\mathcal{T}$ consists of a set of objects sharing the same identity or, equivalently, the set of edges connecting those nodes. Specifically, we model each object candidate $o_{i}\in\mathcal{O}$ as a node and each association hypothesis among candidates as an edge in an undirected graph $G=(V,E)$. Within this formulation, nodes $u\in V$ could be classified into valid objects if $u\in\mathcal{O}_{v}$. Analogously, edges can be classified into correct or incorrect hypotheses depending on whether they belong to some trajectory or not. This formulation can be generalized to a hierarchical setting by splitting video sequences into non-overlapping subsequences, i.e., disjoint subgraphs, and progressively merging them [12].
@@
Background on graph hierarchies for MOT. Our model consists of a hierarchy of GNNs operating over the graph formulation described above. Intuitively, starting from detection candidates, at each hierarchy level, GNN hierarchy subsequently merge tracklets from the previous level into longer ones. Nodes and edges are represented by initial embeddings and message passing is used to share information within the graph in order to update node and edge embeddings with richer information. The GNNs are trained to classify edges into active and inactive association hypotheses and they share the same features, network architecture and learnable weights across hierarchy levels. In addition to the edge features reported in  [12], for each object candidate $o_{i}$ we concatenate its timestamp $t_{i}$, normalized bounding box coordinates and dimensions $(x_{i},y_{i},h_{i},w_{i})$ and corresponding confidence from the detector $c_{i}$ and embed it as node features.
Figure 2: Overview of the SPAM training and annotation pipeline. (a) Initial model training on synthetic data. (b) Application of SPAM to generate pseudo-labels without incurring manual annotation costs on a real dataset, followed by self-training on pseudo-labels. (c) Real dataset labeling using pseudo-labels and an uncertainty-based active learning approach.
@@
Classifying nodes. We aim to minimize the annotation effort; therefore, we ask the annotator to operate on a fixed graph of detection candidates, i.e. annotator does not label new detections after the initial set given by detector. This is why we aim to create an initial graph that is over-complete. In order to do so, we use a low confidence threshold for our detector and obtain a large number of detection candidates. While this ensures a high detection recall, it also introduces a larger number of false positive detections that our model needs to address. As we show in experiments, the original SUSHI architecture [12] cannot handle such false positives, which is why we propose a new layer in the hierarchy, $GNN_{node}$. Intuitively, $GNN_{node}$ leverages temporal information to validate object candidates. We use its output node embeddings and feed them to a lightweight classifier to enable false positive detection filtering. By doing so, we can consider a large number of input proposals achieving high detection recall by leveraging the spatio-temporal reasoning ability of our GNN for detection candidates.
@@ 3.3 How to Label on Graph Hierarchies?
Figure 3: Our graph-based labeling pipeline begins with the selection of nodes for annotation. For each node to be annotated, the annotator could be asked to validate the detection, improve the localization by refining the box or perform association.
Graph-based labeling. Utilizing a graph formulation for detection and tracking enables explicit modeling of objects in the scene and their interactions. In this framework, each node of the graph represents an object, allowing a manual annotation of an object to be framed as selecting a node on this graph. Given a set of nodes to be annotated, possible annotator decisions are: (i) accepting/rejecting a detection as we are working on a large set of detection candidates (one click), (ii) refining the bounding-box to improve localization (two clicks, one for the top-left corner and one for the bottom-right corner of the bounding box) and (iii) associating a node across time (one click) as shown in Fig. 3.
Annotation acquisition. The core idea of our approach is to use our model decisions to guide the annotation process. For the vast majority of scenarios, we directly use our model’s predictions as pseudo-labels. For a small subset of more challenging decisions where our model exhibits uncertainty, we validate our decisions with an annotator as described above. Within our graphs, we adopt an uncertainty-based approach to identify which nodes require annotations. Specifically, for each node $v\in V$, we compute its uncertainty as $\text{uncert}(v)\vcentcolon=\text{max}_{u\in\text{N}_{v}}H({\hat{y}_{(v,u)}})$, where $\text{N}_{v}$ denotes the set of neighboring nodes of $v$ and $\hat{y}_{(v,u)}$ is the edge prediction for edge $(v,u)$:
@@
At the $GNN_{node}$ level, where we have node predictions instead of edge predictions, we simply compute $\text{uncert}(v)$ as $H({\hat{y}_{(v)}})$ instead. Subsequently, we refer the nodes with the highest uncertainty scores to the annotator for manual annotation and use the remaining predictions as pseudo-labels for the rest of the nodes.